USE THE FOLLOWING INFORMATION TO ANSWER THE NEXT (6) QUESTIONS:
The following is a December 31, 2019, post-closing trial balance for Desert Company:

Account Title Debits Credits
Cash $ 83,000
Accounts Receivable (net of Allowance) $280,000
Prepaid Expenses $ 32,000
Investments $ 65,000
Land $175,000
Buildings (net) $160,000
Equipment (net) $145,000
Accounts Payable $ 73,000
Accrued Expenses $ 45,000
Unearned Revenue $150,000
Notes Payable $300,000
Common Stock $200,000
Retained Earnings $172,000
     TOTALS $940,000 $940,000



Additional Information:

1. The cash account includes $15,000 set aside in a legally restricted fund to pay bonds payable that mature in 2027 and a $13,000 cash surrender value of a life insurance policy on the company's CEO.

2. The accounts receivable balance consists of the following:
      a. Amounts owed by customers with debit balances $ 235,000
      b. Customer accounts with credit balances 10,000
      c. Allowance for uncollectible accounts - trade customers (15,000)
      d. Non-trade note receivable due in three years 65,000
      e. Interest receivable on note due in four months     5,000
  Total $280,000

3. The land account includes land which cost $75,000 that the company purchased for speculative purposes and is currently held for sale. The remaining $100,000 is the cost of land on which the company's office building resides. The equipment account includes idle machinery with a book value of $45,000.

4. The investments account is classified as Available for Sale Securities and includes an investment of $25,000 in bonds that mature July 1, 2020. Of the remaining investments balance, management intends to hold for at least the next three years. All investments in the portfolio have already been marked-to-market and are reported at Fair Value.

5. The unearned revenue represents customer prepayments for magazine subscriptions. Subscriptions are for three years and will be earned evenly over each of the years beginning January 1, 2020.

6. The notes payable account consists of the following:
          a. a $50,000 note due in six months.
          b. a $100,000 bond due in six years.
          c. a $150,000 note due in three annual installments of $50,000 each, with the next installment due Nov. 1, 2020.
          *Interest on all notes has been properly accrued and is included in accrued expenses.



Required:

Using this information, determine the amount of Current Assets, after all adjustments and corrections have been made

Using the information in #1 above, determine Total PP&E (net) after all adjustments and corrections have been made:

Using the information in #1 above, determine Total Other Assets after all adjustments and corrections have been made:

Using the information in #1 above, determine Total Long-Term Liabilities after all corrections have been made:
Answer #1

Answer 1 - Total Amount of Current Assets (After Adjustments)

Cash - Value in trial balance = 83,000

Less: Legally restricted funds to pay bonds = 15,000

Less: Cash surrender value of Life insurance policy = 13,000

Adjusted value of cash = $ 55,000

Account Receivable - Value in trial balance = 2,80,000

Add : Customer account with credit balance = 10,000

Less : Non Trade note receivable =  65,000

  Adjusted value of Account Receivable =$ 2,25,000

So total value of current assets after adjustment = Cash + Account Receivable + Prepaid Expenses + Value of Bonds (Maturing July 1,2020)

= 55,000 + 2,25,000 + 32,000 + 25,000

= $ 3,37,000

Answer 2

Total value of long term investment after adjustment = Value given in trial balance - Value of bonds (Maturing on July 1, 2020)

= 65,000 - 25,000

  = $ 40,000

Answer 3 Total Value of PPE (After adjustment)

Land - Value given in balance sheet - Land held for sale

= 1,75,000 - 75,000

= $ 1,00,000

Total value of PPE = Land + Equipment + Building

  = 1,00,000 + 1,45,000 + 1,60,000

= $ 4,05,000

Answer 4 - Value of other assets

Other Assets = Land held for sale (Purchased for speculative purpose)

= $ 75,000

Answer 5- Total value of current liabilities (After Adjustments)

Current liabilities = Unearned revenue + Accounts payable + Accrued expenses + Notes Payable

Unearned Revenue = 1,50,000 - Unearned revenue for 2 years

= 1,50,000 - 1,00,000

= $ 50,000

Notes Payable = 3,00,000 - Bonds due in next 6 years - 2 Annual installments of notes due

= 3,00,000 - 1,00,000 - (50,000 * 2)

= $ 1,00,000

So Current liabilities = 50,000 + 1,00,000 + 45,000 + 73,000

= $ 2,68,000

Answer 6- Value of Long term liabilities -

Long term liabilities = Bonds due in 6 years + 2 Annual installment of notes due + Unearned revenue for 2 years

= 1,00,000 + 1,00,000 + 1,00,000

= $ 3,00,000